AM3 -Esercitazione 7 A.A. 2003-2004

Prof. Luigi Chierchia, Dott. Laura Di Gregorio 23 aprile 2004

Esercizio 1

Calcolare

∫∫_{D} \frac{sin y^{2}}{y} dydx

dove

D = \{(x, y) ∈ ℝ^{2} : 0 ≤ x ≤ y^{2}, 0 < y ≤ \sqrt{π}\}.

Esercizio 2

Calcolare

∫∫_{x^{2}+y^{2}<1}(x − y) cos(x^{20} + y^{20}) dxdy.

Esercizio 3

Calcolare

∫_{0}^{π} ∫_{0}^{1}(x sin y − x^{2}y) dx dy

e verificare che vale il teorema di Fubini.

Esercizio 4

Calcolare i seguenti integrali:

a) ∫∫_{D} \frac{x^{2}}{y^{2}} dx dy dove D = \{(x, y) ∈ ℝ^{2} : 1 ≤ x ≤ 2, \frac{1}{x} ≤ y ≤ x\}

b) ∫∫_{D} y^{3}e^{x} dx dy dove D = {(x, y) ∈ ℝ^{2} : y ≥ 0, x ≤ 1, x ≥ y^{2}} c) ∫∫_{D} xy dx dy dove D = {(x, y) ∈ ℝ^{2} : x + y ≥ 1, x^{2} + y^{2} ≤ 1}